OpenStudy (anonymous):

A square has the same area as a circle of diameter 4. What is the length of each side of the square?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the answer was so obvious

OpenStudy (anonymous):

2 and the square root of pi times 2 and the square root of pi

OpenStudy (cwrw238):

area of the circle = pi*2^2 = 4pi side of square = sort(4pi) = 2 sort pi

OpenStudy (anonymous):

ty, master kitty cat

OpenStudy (cwrw238):

yw big bird

OpenStudy (cwrw238):

- she's female